# Squeezed Diffusion Models

### v1

@@ Abstract
Diffusion models typically inject isotropic Gaussian noise, disregarding structure in the data. Motivated by the way quantum squeezed states redistribute uncertainty according to the Heisenberg uncertainty principle, we introduce Squeezed Diffusion Models (SDM), which scale noise anisotropically along the principal component of the training distribution. As squeezing enhances the signal-to-noise ratio in physics, we hypothesize that scaling noise in a data-dependent manner can better assist diffusion models in learning important data features. We study two configurations: (i) a Heisenberg diffusion model that compensates the scaling on the principal axis with inverse scaling on orthogonal directions and (ii) a standard SDM variant that scales only the principal axis. Counterintuitively, on CIFAR-10/100 and CelebA-64, mild antisqueezing - i.e. increasing variance on the principal axis - consistently improves FID by up to 15% and shifts the precision–recall frontier toward higher recall. Our results demonstrate that simple, data-aware noise shaping can deliver robust generative gains without architectural changes.
@@ 1 Introduction
Score-based diffusion models have become a standard tool for high-fidelity image generation, achieving state-of-the-art performance on datasets from CIFAR-10 to ImageNet (Ho et al., 2020; Dhariwal and Nichol, 2021). While diffusion models typically rely on adding standard isotropic Gaussian noise during the training process, there is evidence that this standard "one-size-fits-all" approach can degrade generative quality in certain contexts like in natural images which have an anisotropic power distribution in frequency space (Falck et al., 2025). This has motivated the investigation of methods to engineer the noise used during the diffusion process in a data-dependent way to overcome these shortcomings (Huang et al., 2024; Sahoo et al., 2024).
The field of quantum metrology Clerk et al. (2008), has developed an extensive array of techniques to manipulate noise in the context of precision measurement. One such technique is quantum squeezing (Walls, 1983), which redistributes noise so that variance is reduced along a “measurement” axis of interest, boosting signal-to-noise ratio (SNR) and enabling weak-signal extraction. Inspired by this, we introduce Squeezed Diffusion Models (SDM), which anisotropically scale noise along the dataset’s principal component. We hypothesize this encourages the model to learn more semantically meaningful patterns during training, improving sample quality in generation.
To test this, we investigated two model variants: (1) a Heisenberg diffusion model that scales the principal direction and counter-scales the orthogonal ones and (2) a "standard" SDM that only scales the principal direction. Surprisingly, we observed that best performance came by mildly antisqueezing along the principal axis - i.e. injecting noise and degrading SNR - across multiple datasets.
@@
by reducing variance in one physical variable $\Delta A$ while enlarging it in the conjugate one $\Delta B$ but keeping the product constant. Squeezed states boost signal-to-noise along a chosen axis and underpin precision measurements such as gravitational-wave Ganapathy et al. (2023) and dark matter Backes et al. (2021) detectors.
@@
Recent work has showed that alternative noise schedules can impact generative performance Chen (2023). Improvements have been demonstrated with for example the cosine schedule Dhariwal and Nichol (2021) and variationally optimized schedules Kingma et al. (2021). Other work has focused on changing noise structure in the frequency domain by e.g. equalizing or biasing certain frequencies Falck et al. (2025); Huang et al. (2024); Hao Phung (2023). Closer to our aim, other work has demonstrated benefits by learning a multivariate data-dependent noise process Sahoo et al. (2024). Our approach differs by using a simple, physics-inspired PCA alignment to apply a single-parameter anisotropic scaling.
@@ 2 Method: Squeezed Diffusion
As in a standard DDPM, an SDM corrupts a clean datapoint $x_{0}$ with additive Gaussian noise over a schedule $\{\beta_{t}\}_{t=1}^{T}\subset(0,1)$. In an SDM however, at each step we anisotropically scale the noise with a squeeze matrix $S_{t}(s)$ whose strength is controlled by a single hyper-parameter $s$ and which may also vary with $t$:
@@
which is still Gaussian with mean $\sqrt{\bar{\alpha}_{t}}\,x_{0}$ and covariance $\Sigma_{t}=\sum_{i=1}^{t}\bigl{(}(1-\alpha_{i})\prod_{j=i+1}^{t}{\alpha_{j}}\bigr{)}S_{i}S_{i}^{\!\top}.$
@@
With this change of variable, eq.˜2 becomes approximately equal to a standard DDPM in $\tilde{x}_{t}$, which allows us to reuse the Ho et al. (2020) posterior:
@@
where $\tilde{\beta}_{t}=(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})\,\beta_{t}$. Finally we re-squeeze with the next matrix $S_{t-1}$ to stay in squeezed coordinates:
@@
To choose the squeezing direction, we do a PCA decomposition of the data in RGB space at the individual pixel level. For large natural image sets, this aligns with the opponent-colour axes luminance, red–green, and blue–yellow Buchsbaum and Gottschalk (1983); Ruderman et al. (1998). Let $\hat{v}$ be the unit vector in the direction of the first principal component, which for RGB images is the luminance. We can define the squeeze matrices $S_{t}^{\mathrm{HDM}}$ for the Heisenberg diffusion model and $S_{t}^{\mathrm{SDM}}$ for the standard squeezed diffusion model:
@@
Here $s$ is a hyperparameter corresponding to the squeezing strength and $n$ is the dimension of the vector space on which squeezing acts. For the case of RGB pixels, $n=3$. Note that $S_{t}^{\mathrm{SDM}}$ is derived from $S_{t}^{\mathrm{HDM}}$ by just discarding the antisqueezing in the subspace orthogonal to $\hat{v}\hat{v}^{\top}$. These matrices mirror the exponential form of electromagnetic squeezing operators in quantum mechanics Gerry and Knight (2005):
$S(z)=\exp\left(\frac{z^{*}\hat{a}^{2}+z(\hat{a}^{\dagger})^{2}}{2}\right)$ (10)
For $z=re^{i\theta}$, $S(z)$ scales the noise by $e^{-r}$ in the $\theta$ direction in the phase space determined by the electromagnetic creation and annihilation operators $\hat{a}^{\dagger}$, $\hat{a}$, just as the $S_{t}$ do so by $e^{-s}$ in the direction determined by $\hat{v}$.
@@ 3 Results
Our experiments consist of ablation studies to evaluate squeezed diffusion over a variety of standard image datasets. The results are shown in fig.˜2.
@@
On CIFAR-10 Krizhevsky (2009), FID-10K shows a consistent U-shaped dependence on squeeze strength (fig.˜2(a)) across multiple random seeds. The average baseline DDPM FID was 39.8 while the SDM reached an FID of 33.6 at $s\approx-0.4$ - a $\sim$15% improvement. Notably, this is in the antisqueezing regime with degraded SNR.
Next, with a fixed seed, we compare the standard squeezed diffusion model against the uncertainty preserving Heisenberg variant on CIFAR-10. Both the FID-10K (fig.˜2(b)) and Inception Score (fig.˜2(c)) follow the same U-shaped trend in $s$, with optima near $s\approx-0.4$ (FID) and $s\approx-0.3$ (IS). SDM outperforms HDM in the mild antisqueezing regime.
To check the robustness of squeezing, we evaluated SDM generative performance on CIFAR-100 Krizhevsky (2009) in fig.˜2(d) and CelebA-64 Liu et al. (2015) in fig.˜2(e). The parabolic shape is reproduced across datasets, with the optimum achieved in the mild antisqueezing regime again. The exact location of the optimum varies across datasets, with best performance on CIFAR-100 at $s\sim-0.3$ and on CelebA-64 at $s\sim-0.2$.
To gain further insight into the impact of squeezing vs antisqueezing, we analyze the precision and recall metrics of our models on CIFAR-10 Sajjadi et al. (2018); Kynkäänniemi et al. (2019); Obukhov et al. (2020). Figure˜2(f) shows that squeezing the SDM noise corresponds to reducing recall while getting marginal gains in precision. By contrast antisqueezing the noise produces large gains in recall at minor cost in precision up to a point. Overall generative quality, as measured by F-score, once again improves up to a point showing parabolic behavior.
@@ 4 Conclusion and Future Work
Across three image datasets we observe a stable U-shaped dependence of sample quality on the squeeze strength: modest antisqueezing enlarges variance along luminance, boosting recall at little precision cost, whereas strong squeezing or antisqueezing harms both. The standard SDM variant matches or outperforms the Heisenberg counterpart, underscoring that diffusion models need not strictly conserve uncertainty to improve generative performance. Future work could include scaling to higher resolution datasets and exploring frequency-dependent or modality-specific squeezing such as audio. Overall, SDM offers a principled method to manipulate noise that delivers robust gains in generative quality.
@@ A.1 Implementation
We use the HuggingFace diffusers library von Platen et al. with a UNet2DModel backbone. To apply squeezing, we keep the UNet architecture intact and subclass the DDPMScheduler class overriding the add_noise and step methods. All other training and sampling code remains standard. The model uses a linear noise schedule. Training is done with 1000 timesteps while sampling is done using 50. We also used EMA with a value of EMA decay set to $0.9999$. All experiments were run on a single NVIDIA A100 (40GB) GPU with mixed precision (fp16) via accelerate. Code for these experiments is available at https://github.com/joe-singh/squeezing.

### v2

@@ Abstract
Diffusion models typically inject isotropic Gaussian noise, disregarding structure in the data. Motivated by the way quantum squeezed states redistribute uncertainty according to the Heisenberg uncertainty principle, we introduce Squeezed Diffusion Models (SDM), which scale noise anisotropically along the principal component of the training distribution. As squeezing enhances the signal-to-noise ratio in physics, we hypothesize that scaling noise in a data-dependent manner can better assist diffusion models in learning important data features. We study two configurations: (i) a Heisenberg diffusion model that compensates the scaling on the principal axis with inverse scaling on orthogonal directions and (ii) a standard SDM variant that scales only the principal axis. Counterintuitively, on CIFAR-10/100 and CelebA-64, mild antisqueezing – i.e. increasing variance on the principal axis – consistently improves FID by up to 15% and shifts the precision–recall frontier toward higher recall. Our results demonstrate that simple, data‑aware noise shaping can deliver robust generative gains without architectural changes.
@@ 1 Introduction
Score-based diffusion models have become a standard tool for high-fidelity image generation, achieving state-of-the-art performance on datasets from CIFAR-10 to ImageNet [ho2020denoising, dhariwal2021diffusion]. While diffusion models typically rely on adding standard isotropic Gaussian noise during the training process, there is evidence that this standard "one-size-fits-all" approach can degrade generative quality in certain contexts like in natural images which have an anisotropic power distribution in frequency space [falck2025fourier]. This has motivated the investigation of methods to engineer the noise used during the diffusion process in a data-dependent way to overcome these shortcomings [huang2024bluenoise, sahoo2024mulan].
The field of quantum metrology [clerk2008quantumnoise], has developed an extensive array of techniques to manipulate noise in the context of precision measurement. One such technique is quantum squeezing [walls1983squeezing], which redistributes noise so that variance is reduced along a “measurement” axis of interest, boosting signal-to-noise ratio (SNR) and enabling weak-signal extraction. Inspired by this, we introduce Squeezed Diffusion Models (SDM), which anisotropically scale noise along the dataset’s principal component. We hypothesize this encourages the model to learn more semantically meaningful patterns during training, improving sample quality in generation.
To test this, we investigated two model variants: (1) a Heisenberg diffusion model that scales the principal direction and counter-scales the orthogonal ones and (2) a "standard" SDM that only scales the principal direction. Surprisingly, we observed that best performance came by mildly antisqueezing along the principal axis – i.e. injecting noise and degrading SNR – across multiple datasets.
@@
by reducing variance in one physical variable $\Delta A$ while enlarging it in the conjugate one $\Delta B$ but keeping the product constant. Squeezed states boost signal-to-noise along a chosen axis and underpin precision measurements such as gravitational-wave [ligo2024squeezing] and dark matter [haystac2021squeezing] detectors.
@@
Recent work has showed that alternative noise schedules can impact generative performance [chen2023importancenoiseschedulingdiffusion]. Improvements have been demonstrated with for example the cosine schedule [dhariwal2021diffusion] and variationally optimized schedules [kingma2023variational]. Other work has focused on changing noise structure in the frequency domain by e.g. equalizing or biasing certain frequencies [falck2025fourier, huang2024bluenoise, hao2024wavelet]. Closer to our aim, other work has demonstrated benefits by learning a multivariate data-dependent noise process [sahoo2024mulan]. Our approach differs by using a simple, physics-inspired PCA alignment to apply a single-parameter anisotropic scaling.
@@ 2 Method: Squeezed Diffusion
As in a standard DDPM, an SDM corrupts a clean datapoint $x_{0}$ with additive Gaussian noise over a schedule $\{\beta_{t}\}_{t=1}^{T}\subset(0,1)$. In an SDM however, at each step we anisotropically scale the noise with a squeeze matrix $S_{t}(s)$ whose strength is controlled by a single hyper‑parameter $s$ and which may also vary with $t$:
@@
As in a standard DDPM, we have that $q(x_{t}|x_{0})\sim\mathcal{N}\left(\sqrt{\bar{\alpha}}_{t}x_{0},\Sigma_{t}\right)$ but instead now the covariance is modified to be $\Sigma_{t}=\sum_{i=1}^{t}\bigl((1-\alpha_{i})\prod_{j=i+1}^{t}{\alpha_{j}}\bigr)S_{i}S_{i}^{\!\top}.$
@@
With this change of variable, eq.˜2 becomes approximately equal to a standard DDPM in $\tilde{x}_{t}$, which allows us to reuse the ho2020denoising posterior:
@@
where $\tilde{\beta}_{t}=(1-\bar{\alpha}_{t-1})/(1-\bar{\alpha}_{t})\,\beta_{t}$. Finally we re‑squeeze with the next matrix $S_{t-1}$ to stay in squeezed coordinates:
@@
To choose the squeezing direction, we do a PCA decomposition of the data in RGB space at the individual pixel level. For large natural image sets, this aligns with the opponent‑colour axes luminance, red–green, and blue–yellow [buchsbaum1983trichromacy, Ruderman:98]. Let $\hat{v}$ be the unit vector in the direction of the first principal component, which for RGB images is the luminance. We can define the squeeze matrices $S_{t}^{\mathrm{HDM}}$ for the Heisenberg diffusion model and $S_{t}^{\mathrm{SDM}}$ for the standard squeezed diffusion model:
@@
Here $s$ is a hyperparameter corresponding to the squeezing strength and $n$ is the dimension of the vector space on which squeezing acts. For the case of RGB pixels, $n=3$. Note that $S_{t}^{\mathrm{SDM}}$ is derived from $S_{t}^{\mathrm{HDM}}$ by just discarding the antisqueezing in the subspace orthogonal to $\hat{v}\hat{v}^{\top}$. These matrices mirror the exponential form of electromagnetic squeezing operators in quantum mechanics [quantumoptics2005]:
$S(z)=\exp\left(\frac{z^{*}\hat{a}^{2}-z(\hat{a}^{\dagger})^{2}}{2}\right)$ (10)
For $z=re^{i\theta}$, $S(z)$ scales the noise by $e^{-r}$ in the $\theta$ direction in the phase space determined by the electromagnetic creation and annihilation operators $\hat{a}^{\dagger}$, $\hat{a}$, just as the $S_{t}$ do so by $e^{-s}$ in the direction determined by $\hat{v}$. Note also that with this parametrisation, there is a smooth, continuous path between an SDM and an isotropic DDPM (achieved at $s=0$). This mirrors the way squeezing of vacuum noise is inherently continuous, and it situates SDM as a principled middle ground between isotropic, data-unaware noise and fully data-aware approaches such as covariance-matching diffusion [sahoo2024mulan].
@@ 3 Results
Our experiments consist of ablation studies to evaluate squeezed diffusion over a variety of standard image datasets. All experiments were run on a single NVIDIA A100 (40GB) GPU with mixed precision (fp16) via accelerate. The results are shown in fig.˜2.
@@
On CIFAR-10 [krizhevsky2009learning], FID-10K shows a consistent U-shaped dependence on squeeze strength (fig.˜2(a)) across multiple random seeds. The average baseline DDPM FID was 39.8 while the SDM reached an FID of 33.6 at $s\approx-0.4$ – a $\sim$15% improvement. Notably, this is in the antisqueezing regime with degraded SNR.
Next, with a fixed seed, we compare the standard squeezed diffusion model against the uncertainty preserving Heisenberg variant on CIFAR‑10. Both the FID-10K (fig.˜2(b)) and Inception Score (fig.˜2(c)) follow the same U‑shaped trend in $s$, with optima near $s\approx-0.4$ (FID) and $s\approx-0.3$ (IS). SDM outperforms HDM in the mild antisqueezing regime.
To check the robustness of squeezing, we evaluated SDM generative performance on CIFAR-100 [krizhevsky2009learning] in fig.˜2(d) and CelebA-64 [liu2015faceattributes] in fig.˜2(e). The parabolic shape is reproduced across datasets, with the optimum achieved in the mild antisqueezing regime again. The exact location of the optimum varies across datasets, with best performance on CIFAR-100 at $s\sim-0.3$ and on CelebA-64 at $s\sim-0.2$.
To gain further insight into the impact of squeezing vs antisqueezing, we analyze the precision and recall metrics of our models on CIFAR-10 [sajjadi2018assessing, kynkaanniemi2019improved, obukhov2020torchfidelity]. Figure˜2(f) shows that squeezing the SDM noise corresponds to reducing recall while getting marginal gains in precision. By contrast antisqueezing the noise produces large gains in recall at minor cost in precision up to a point. Overall generative quality, as measured by F-score, once again improves up to a point showing parabolic behavior.
@@ 4 Conclusion and Future Work
Across three image datasets we observe a stable U‑shaped dependence of sample quality on the squeeze strength: modest antisqueezing enlarges variance along luminance, boosting recall at little precision cost, whereas strong squeezing or antisqueezing harms both. The standard SDM variant matches or outperforms the Heisenberg counterpart, underscoring that diffusion models need not strictly conserve uncertainty to improve generative performance. Future work could include scaling to higher resolution datasets and exploring frequency-dependent or modality-specific squeezing such as audio. Overall, SDM offers a principled method to manipulate noise that delivers robust gains in generative quality.
@@ A.1 Implementation
We use the HuggingFace diffusers library [von_Platen_Diffusers_State-of-the-art_diffusion] with a UNet2DModel backbone. To apply squeezing, we keep the UNet architecture intact and subclass the DDPMScheduler class overriding the add_noise and step methods. All other training and sampling code remains standard. The model uses a linear noise schedule. Training is done with 1000 timesteps while sampling is done using 50. We also used EMA with a value of EMA decay set to $0.9999$. Code for these experiments is available on Github at https://github.com/joe-singh/squeezing.